For the 900 trading days from January 2003 through July 2006, the daily closing price of IBM stock (in$) is well modeled by a Normal model with mean $85.60 and standard deviation $6.20. According to this model, what is the probability that on a randomly selected day in this period the stock price closed above $91.80?
Answer by stanbon(75887)

For the 900 trading days from January 2003 through July 2006, the daily closing price of IBM stock (in$) is well modeled by a Normal model with mean $85.60 and standard deviation $6.20. According to this model, what is the probability that on a randomly selected day in this period the stock price closed above $91.80?
-------------------
z(91.80) = (91.80-85.60)/6.2 = 1
---
P(x > 91.80) = P(z > 1) = 0.1587
================
Cheers,
Stan H.
================